OpenStudy (anonymous):

What is the vertex of the graph of y = |2x - 6| - 4? A. (3, -4) B. (-3, -4) C. (2, -4) D. (6, -4)

OpenStudy (anonymous):

By definition of |x|, |2x-6|-4 = {2x-6-4 = 2x-10 if 2x-6≥0 {-(2x-6)-4 = 2-2x if 2x-6<0 So , the graph is composed of two lines, one with positive slope and other with negative slope. So the vertex must be the instersection point between those lines. So the x coordinate of that point, is given by the equation 2x-10 = 2-2x Solving for x and y, the answer happens to be (3,-4)